OpenStudy (anonymous):

plase solve and explain the question ill post to down :

OpenStudy (anonymous):

\[\int\limits_{}\sqrt{x}\sqrt{a^{2}x^{2}}dx\]

OpenStudy (amistre64):

ax.sqrt(x) dx; u = sqrt(x); u^2 = x; 2u du = dx ..... maybe a.u^2.u.2u.du

OpenStudy (amistre64):

{S} 2a.u^4.du ; the 2a acts like a constant perhaps? \[\frac{2au^5}{5}\]

OpenStudy (amistre64):

u^5 = u^2.u^2.u = x.x.sqrt(x) \[\frac{2a}{5}x^2\sqrt{x}+C\] maybe

OpenStudy (anonymous):

maybe :)

OpenStudy (amistre64):

it works lol

OpenStudy (anonymous):

u took u = sqrt(x) so that it been u^2*a*u ^^ then its u^3*adx

OpenStudy (amistre64):

\[\frac{2a}{5}*(2x\sqrt{x}+\frac{x^2}{2\sqrt{x}})\]

OpenStudy (anonymous):

why isnt \[\sqrt{x}\times \sqrt{x^3}=x^2\]?

OpenStudy (anonymous):

and du =u^2 so its now u^3.a.u^2 then its u^5.a / 5 + c

OpenStudy (anonymous):

if the problem says (my eyes not what they ought to be) \[\int\sqrt{x}\sqrt{a^2x^3}dx\] then isnt this the same as \[a\int x^2 dx\]?

OpenStudy (anonymous):

just get \[\frac{ax^3}{3}\] and be done

OpenStudy (anonymous):

nono its X^2 not X^3 in this sqrt :D

OpenStudy (amistre64):

\[\frac{2a}{5}(\frac{4x^2+x^2}{2\sqrt{x}})=\frac{2a}{5}(\frac{5x^2}{2\sqrt{x}})\] \[2a(\frac{x^2\sqrt{x}}{2x})=a(x\sqrt{x})=ax\sqrt{x}\]

OpenStudy (anonymous):

ooooooooooooooh sorry

OpenStudy (anonymous):

i think its \[a*u^{5}/5\]

OpenStudy (anonymous):

and + C

OpenStudy (amistre64):

its: 2a x^5 ------ +C :) 5

OpenStudy (anonymous):

then i believe you get \[\frac{2a}{5}\sqrt{x^5}\]

OpenStudy (anonymous):

amistre root over the \[x^5\]

OpenStudy (anonymous):

2 ?? where this 2 came from ?? i thought that dx was du^2

OpenStudy (anonymous):

forget u and du etc

OpenStudy (anonymous):

just rewrite w/ exponents and use power rule

OpenStudy (anonymous):

\[\sqrt{x}\sqrt{a^2x^2}=\sqrt{a^2}x^{\frac{3}{2}}\]

OpenStudy (anonymous):

hmm , why u don't get a^2 out of the swrt?

OpenStudy (anonymous):

hmm , why u don't get a^2 out of the swrt?

OpenStudy (anonymous):

because it is always possible that a < 0

OpenStudy (anonymous):

and i don't want to fret about rewriting as |a|

OpenStudy (anonymous):

wel but ^2 always will equal to +

OpenStudy (anonymous):

so i left it in the redical

OpenStudy (anonymous):

exactly. which is why i cannot pull it out

OpenStudy (anonymous):

because if i write it as \[a x^{\frac{3}{2}}\] this would be incorrect if say a = -3

OpenStudy (anonymous):

satalllite i think its easier to define ıu and du :D

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

just have \[\int x^{\frac{3}{2}}dx=\frac{2}{5}x^{\frac{5}{2}}\] finish

OpenStudy (anonymous):

well its goo dtoo :D

OpenStudy (anonymous):

just use power rule that is all. but of course you are free to use whatever

OpenStudy (anonymous):

thanks :D